OpenStudy (anonymous):

show all you work the length of a football field is 67 yards greater than the width. the perimeter of the field is 346 yards. find the length and width of the football field?

OpenStudy (anonymous):

 4x+134=345 -134 -134 4x+212 /4 /4 x=53 length= 120 width=53